OpenStudy (anonymous):

A block is projected along a frictionless inclined plane at an angle of 32 degrees with an initial speed of 3.5m/s. How far up the plane does the block go?

OpenStudy (anonymous):



OpenStudy (anonymous):

could this be the diagram to represent the whole story?

OpenStudy (potatoes.ramu):

Can you use tan32 to find the slope?

OpenStudy (potatoes.ramu):

That way you have the total distance right?

OpenStudy (potatoes.ramu):

And you know the accleration of the block is mg

OpenStudy (potatoes.ramu):

Then... you could probably use the formula v^2=u^2+2as?

OpenStudy (anonymous):

ya, I said in my working gsine32 = a= \[9.8x0.59=5.19\] after that used \[v ^{2} = u ^{2}-2as= 3.5^{2}-2x5.19 s.\]

OpenStudy (potatoes.ramu):

yay! So our thinking was along the same lines.. again :P But what do you do after that?

OpenStudy (anonymous):

\[s = \frac{ -3.5^{2} }{ -5.19 } = 2.36\]

OpenStudy (potatoes.ramu):

Is that the answer?

OpenStudy (anonymous):

This is the one I came up with, I think it go as much as 2.36m???

OpenStudy (potatoes.ramu):

I didn't understand how you found the time...

OpenStudy (anonymous):

oh, S is for displacement in the text that I am using, sorry about that man

OpenStudy (potatoes.ramu):

Woman* Oh right.. I thought it was seconds.. oops

OpenStudy (potatoes.ramu):

The Free body diagram looks right.. Why'd you ask this question when you already know the answer though? :P

OpenStudy (anonymous):

I am sorry I left wo- before man ha ha ha. I am revising alone, so with physics and mathematics you always underline a certain answer , even if it is wrong. So i have to ask. I think u understand what I am trying to say. And thanx for helping me.

OpenStudy (potatoes.ramu):

I have no idea how I helped, but you're welcome :) Message if you have any more physics questions! They help me too!!

OpenStudy (anonymous):

The fact that you were understanding my approach , and the direction of my working( of cause except for the s part), means that I was presenting the whole thing well. Let me say u were supporting?, Anyway, I will always mention you when I have a question, actually, let me close this one , and write the other one which I wasn't so sure of.